There is 355 students altogether. The school has 25 more females than males how many female and males does the school have.
Answer by stanbon(75887)

There is 355 students altogether. The school has 25 more females than males how many female and males does the school have.
------
Equations:
f + m = 355
f = m+25
-------
Substitute for m and solve for "f":
f + f-25 = 355
-----
2f = 380
f = 190 (# of females)
===============
Cheers,
Stan H.
===============